# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@ 5.2.2 Zero-dimensional critical complete intersections
Assume that $D\subset(\mathbb{Z}^{n})^{\star}$ is a hyperplane. It is uniquely determined by the common zero $D^{\star}\in\mathbb{Q}\mathbb{P}^{n}$ of all affine linear functions $\alpha\in D$. We also denote by $Q^{\star},Q\in\mathbb{Q}\mathbb{P}^{n}$ the hyperplane in $(\mathbb{Z}^{n})^{\star}$ consisting of all affine functions vanishing at $Q$. Let $\mathfrak{A}_{\mathbf{A}}(D^{\star})$ be the set $\{\Gamma\,|\,\Gamma$ is a face of the convex hull of $\mathbf{A}$, and the affine span of $\Gamma$ contains $Q\}$.
@@
[Es17, Section 2.7.] Consider a generic polynomial $f$ with support $\mathbf{A}$. Critical complete intersection defined by $f$ is indeed a complete intersection. The system $\partial_{D}(f)$ is degenerate at a face $\Gamma\subset\operatorname{Conv}(\mathbf{A})$ (in the Bernstein-Kouchnirenko-Khovanskii sense) only if $\Gamma\in\mathfrak{A}_{\mathbf{A}}(D^{\star})$.
@@
(the statement appears in [Es17, Lemma 2.50.], see [Es24b, Theorem 1.8] for a more detailed proof) Consider a finite set $\mathbf{A}\subset\mathbb{Z}^{n}$ and a hyperplane $D$ in $(\mathbb{Z}^{n})^{\star}$, then, for a generic Laurent polynomial $f(x)=\sum_{z\in A}c_{a}z^{a}$, the zero-dimensional critical complete intersection $\{\partial_{D}f=0\}$ consists of $\nu_{D^{\star}}^{e}(\mathbf{A})$ points (counted with multiplicity).
@@
The above lemma is proved in [Es24b, Theorem 1.8] for $D^{\star}=0$. This immediately implies the lemma for arbitrary $D^{\star}\in\mathbb{Z}^{n}$ since it corresponds to the shift of $\mathbf{A}$ by $-D^{\star}$. The proof for arbitrary $D^{\star}\in\mathbb{Q}\mathbb{P}^{n}$ is the same; it can be also proved by constructing an appropriate set $\mathbf{A}^{+1}\subset\mathbb{Z}^{n+1}$, so that $\nu^{e}_{D^{\star}}(\mathbf{A})=\nu^{e}_{O}(\mathbf{A}^{+1})$ (as in Remark 5.42).
@@
In [Es17, Lemma 2.50.] it is said that the formula from Lemma 5.21 ‘‘is the Bernstein formula [Ber75] with some roots hidden at infinity (more precisely, at the orbits of the $\mathbf{A}$-toric variety, corresponding to the faces from $\mathbf{A}$)’’. It is easier to see that if we invert the formula:
@@
The number of critical points #$\#\{\frac{\partial f}{x_{1}}=\dots=\frac{\partial f}{x_{n}}=0\}$ of a generic polynomial $f$ with support $\mathbf{A}$ in the torus equals $\nu^{e}_{O}(\mathbf{A})$, where $O$ is the origin.
@@
5.3 Comparison of Newton numbers: $\nu$, $\nu^{\ell}$ and $\nu^{e}$
@@
Both $\nu^{\ell}$ and $\nu^{e}$ are non-negative (by Theorem 4.6 and Lemma 5.21), while the usual $\nu$ is generally not (see [GKZ94, Example 2.5, p.353]).
@@
Consider a set $\mathbf{P}$ such that $P=\operatorname{Conv}(\mathbf{P})$ is a convenient polytope in the cone $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ and a point $X\in\mathbb{Z}^{n-m}$. Then we have:
@@
Recall that the sections $\partial_{e_{1}}f=\dots=\partial_{e_{n}}f$ of $I_{P}$ have only isolated intersections on $\mathbb{T}^{\Sigma}$ (see Remark 5.23). Thus the global intersection number is the sum of positive local intersection multiplicities. By Remark 5.4 the sum of multiplicities of $\{\partial_{e_{1}}f=\dots=\partial_{e_{n}}f=0\}$ on $\mathbb{T}^{\Sigma}$ is equal to $\operatorname{Vol}_{\mathbb{Z}}(P)$. The inclusion-exclusion principle (similar to [Oka79, Section 4]) implies that:
$\nu_{C}(P)=mult(D_{f,1}\cdot{\dots}\cdot D_{f,n}\cdot\mathbb{T}^{\Sigma})$
@@
The following remark is implied by the proof of Lemma 5.27 in case of convenient polytopes, and it is also implied by [Sel24, Theorem 5.10 (2)].
@@
The Newton number in the cone $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ is monotonous, i.e., if $P_{1}\subset P_{2}\subset C$ are convenient, then $\nu_{C}(P_{1})\leq\nu_{C}(P_{2})$.
@@
Consider a cone $C$ with boolean face poset $\mathcal{C}$, i.e., its projection along the maximal affine subspace of $C$ is a simplicial cone. Denote its facets by $E_{1},\dots,E_{m}$. Consider a lattice set $\mathbf{P}\subset C$, the polytope $P=\operatorname{Conv}(\mathbf{P})$ and a point $Q\in E_{1}\cap\dots\cap E_{m}$. Suppose that the affine span of any face in $\mathfrak{A}_{\mathbf{P}}(Q)$ is a face of $C$. Then we have the following formula:
$\nu_{Q}^{e}(\mathbf{P})=MV(\operatorname{Conv}(\mathbf{P}\setminus E_{1}),\dots,\operatorname{Conv}(\mathbf{P}\setminus E_{m}),P,\dots,P)$
@@
The conditions of the above lemma are applied, for instance, if $P$ is a convenient and $Q\in E_{1}\cap\dots\cap E_{m}$ is a generic point.
@@
Consider the covectors $\xi_{1},\dots,\xi_{m}$ orthogonal to the facets of $C$. Consider $n-m$ affine functions $\psi_{m+1},\dots,\psi_{n}$ vanishing on $Q$ so that $\xi_{1},\dots,\xi_{m},\psi_{m+1},\dots,\psi_{n}$ form a basis in $Q^{\star}$. The critical complete intersection is defined by the system $\partial_{\xi_{1}}f=\dots=\partial_{\xi_{m}}f=\partial_{\psi_{m}+1}f=\dots\partial_{\psi_{n}}f$. Note that the Newton polytopes of the corresponding polynomials are contained in $Conv(\mathbf{P}\setminus E_{1}),\dots,Conv(\mathbf{P}\setminus E_{m}),P,\dots,P$ respectively. Proposition 5.18 implies that the corresponding system is non-degenerate in the sense of the Bernstein-Kouchnirenko theorem 5.3. This implies that the number of solutions equals the mixed volume. ∎
@@
Consider a finite lattice set $\mathbf{P}$, its convex hull $P=\operatorname{Conv}(\mathbf{P})$, and a point $Q\in\mathbb{Q}^{n}$. Suppose that for any face $F\in\mathfrak{A}_{\mathbf{A}}(Q)$, the set $\mathbf{P}$ is smooth at $\mathbf{P}\cap F$. Then we have $\nu_{Q}^{e}(\mathbf{A})=\nu_{{C_{Q}(\mathbf{P})}}(P)$.

### v2

@@ 5.2.2 Zero-dimensional critical complete intersections
Assume that $D\subset(\mathbb{Z}^{n})^{\star}$ is a hyperplane. It is uniquely determined by the common zero $D^{\star}\in\mathbb{Q}\mathbb{P}^{n}$ of all affine linear functions $\alpha\in D$. We also denote by $Q^{\star},Q\in\mathbb{QP}^{n}$ the hyperplane in $(\mathbb{Z}^{n})^{\star}$ consisting of all affine functions vanishing at $Q$. Let $\mathfrak{A}_{\mathbf{A}}(D^{\star})$ be the set of faces $\Gamma$ of $\operatorname{Conv}(\mathbf{A})$ such that the affine span of $\Gamma$ contains $Q$.
@@
[Es18, Proposition 2.43 and 2.47] Consider a generic polynomial $f$ with support $\mathbf{A}$. The critical complete intersection defined by $\{\partial_{D}(f)=0\}\subset(\mathbb{C}\setminus 0)^{n}$ is indeed a complete intersection (i.e. its dimension is zero). The system $\partial_{D}(f)$ can be degenerate at a face $\Gamma\subset\operatorname{Conv}(\mathbf{A})$ (in the Bernstein-Kouchnirenko sense Theorem 5.4) only if $\Gamma\in\mathfrak{A}_{\mathbf{A}}(D^{\star})$.
@@
(The statement appears in [Es18, Lemma 2.50.], see [Es24b, Theorem 1.8] for a more detailed proof). Consider a finite set $\mathbf{A}\subset\mathbb{Z}^{n}$ and a hyperplane $D$ in $(\mathbb{Z}^{n})^{\star}$, then, for a generic Laurent polynomial $f(x)=\sum_{z\in A}c_{a}z^{a}$, the zero-dimensional critical complete intersection $\{\partial_{D}f=0\}$ consists of $\nu_{D^{\star}}^{e}(\mathbf{A})$ points (counted with multiplicity).
@@
The above lemma is proved in [Es24b, Theorem 1.8] for $D^{\star}=0$. This immediately implies the lemma for arbitrary $D^{\star}\in\mathbb{Z}^{n}$ since it corresponds to the shift of $\mathbf{A}$ by $D^{\star}$. The proof for arbitrary $D^{\star}\in\mathbb{QP}^{n}$ is the same; it can be also proved by constructing an appropriate set $\mathbf{A}^{+1}\subset\mathbb{Z}^{n+1}$, so that $\nu^{e}_{D^{\star}}(\mathbf{A})=\nu^{e}_{O}(\mathbf{A}^{+1})$ (as in Remark 5.49).
@@
In [Es18, Lemma 2.50] it is said that the formula from Lemma 5.22 ‘‘is the Bernstein formula [Ber75] with some roots hidden at infinity (more precisely, at the orbits of the $\mathbf{A}$-toric variety, corresponding to the faces from $\mathbf{A}$)’’. It is easier to see that if we invert the formula:
@@
The number of critical points $\{\frac{\partial f}{\partial x_{1}}=\dots=\frac{\partial f}{\partial x_{n}}=0\}$ of a generic polynomial $f$ with support $\mathbf{A}$ in the torus equals $\nu^{e}_{O}(\mathbf{A})$, where $O$ is the origin.
@@
5.3 Comparison of the Newton numbers: $\nu$, $\nu^{\ell}$ and $\nu^{e}$
@@
Both $\nu^{\ell}$ and $\nu^{e}$ are non-negative (by Theorem 4.6 and Lemma 5.22), while the usual $\nu$ is generally not (see [GKZ94, Example 2.5 on p.353]).
@@
Consider a finite set $\mathbf{P}\subset\mathbb{Z}^{n}$ such that $P=\operatorname{Conv}(\mathbf{P})$ is a convenient polytope in the cone $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ and a point $X\in\mathbb{Z}^{n-m}$. Then we have:
@@
Recall that the sections $\partial_{e_{1}}f=\dots=\partial_{e_{n}}f=0$ of $I_{P}$ have only isolated intersections on $\mathbb{T}^{\Sigma}$ (see Remark 5.24). Thus the global intersection number is the sum of positive local intersection multiplicities. By Remark 5.5, the sum of multiplicities of $\{\partial_{e_{1}}f=\dots=\partial_{e_{n}}f=0\}$ on $\mathbb{T}^{\Sigma}$ is equal to $\operatorname{Vol}_{\mathbb{Z}}(P)$. The inclusion-exclusion principle (similar to [Oka79, Section 4]) implies that:
$\nu_{C}(P)=\operatorname{mult}(D_{f,1}\cdot\ldots\cdot D_{f,n}\cdot\mathbb{T}^{\Sigma})$ (1)
@@
The following remark is implied by the proof of Lemma 5.28 (more precisely, by equation (1)); it is also implied by [Sel24, Theorem 5.10 (2)].
@@
The Newton number in the cone $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ is monotonic, i.e., if $P_{1}\subset P_{2}\subset C$ are convenient, then $\nu_{C}(P_{1})\leq\nu_{C}(P_{2})$.
@@
Consider a cone $C\subset\mathbb{R}^{n},\dim C=n$ with boolean face poset $\mathcal{C}$, i.e., its projection along the maximal affine subspace of $C$ is a simplicial cone. Denote its facets by $E_{1},\dots,E_{m}$. Consider a lattice set $\mathbf{P}\subset C$, the polytope $P=\operatorname{Conv}(\mathbf{P}),\dim P=n$ and a point $Q\in E_{1}\cap\dots\cap E_{m}$. Suppose that the affine span of any face in $\mathfrak{A}_{\mathbf{P}}(Q)$ coincides with the affine span of a face of $C$. Then we have the following formula:
$\nu_{Q}^{e}(\mathbf{P})=\operatorname{MV}(\operatorname{Conv}(\mathbf{P}\setminus E_{1}),\dots,\operatorname{Conv}(\mathbf{P}\setminus E_{m}),\underbrace{P,\dots,P}_{n-m})$
@@
The conditions of the above lemma are applied, for instance, if $P$ is a convenient and $Q\in E_{1}\cap\dots\cap E_{m}$ is a generic point. Also note that the conditions of the lemma imply that if for a face $C^{\prime}\leq C$ its affine span coincides with the affine span of a face in $\mathfrak{A}_{\mathbf{P}}(Q)$, then for any face $C^{\prime}\leq C^{\prime\prime}\leq C$ the affine span of $C^{\prime\prime}$ also coincides with the affine span of a face in $\mathfrak{A}_{\mathbf{P}}(Q)$.
@@
Consider the covectors $\xi_{1},\dots,\xi_{m}$ orthogonal to the facets of $C$. Consider nmn-m affine functions $\psi_{m+1},\dots,\psi_{n}$ vanishing on $Q$ so that $\xi_{1},\dots,\xi_{m},\psi_{m+1},\dots,\psi_{n}$ form a basis in $Q^{\star}$. The critical complete intersection is defined by the system $\partial_{\xi_{1}}f=\dots=\partial_{\xi_{m}}f=\partial_{\psi_{m+1}}f=\dots=\partial_{\psi_{n}}f$. Note that the Newton polytopes of the corresponding polynomials are contained in the following polytopes
$Conv(\mathbf{P}\setminus E_{1}),\dots,Conv(\mathbf{P}\setminus E_{m}),\underbrace{P,\dots,P}_{n-m}$
respectively. Let us verify that this system is non-degenerate in the Bernstein-Kouchnirenko (see Theorem 5.4). This implies that the number of solutions equals the mixed volume.
We need to check that for generic $f$ with support $\mathbf{P}$ for each $\xi\in(\mathbb{R}^{n})^{*}\setminus 0$ the system
$(\partial_{\xi_{1}}f)^{\xi}=\dots=(\partial_{\xi_{m}}f)^{\xi}=(\partial_{\psi_{m+1}}f)^{\xi}=\dots=(\partial_{\psi_{n}}f)^{\xi}=0,$ (2)
considered as system of polynomials with Newton polytopes $Conv(\mathbf{P}\setminus E_{1}),\dots,Conv(\mathbf{P}\setminus E_{m}),P,\dots,P$ respectively, has no solutions in the torus $(\mathbb{C}\setminus 0)^{n}$.
  1. 1.
    Suppose that $P^{\xi}\notin\mathfrak{A}_{\mathbf{P}}(Q)$; then by Proposition 5.19 the system
    $(\partial_{\xi_{1}}f)|_{P^{\xi}}=\dots=(\partial_{\xi_{m}}f)|_{P^{\xi}}=(\partial_{\psi_{m+1}}f)|_{P^{\xi}}=\dots=(\partial_{\psi_{n}}f)|_{P^{\xi}}=0$ (3)
    has no solutions in the torus. This implies that the system (2) has no solutions, since it contains all nonzero polynomials from the latter system (3).
  2. 2.
    Suppose that $P^{\xi}\in\mathfrak{A}_{\mathbf{P}}(Q)$, for instance, the affine span of $P^{\xi}$ coincides with the affine span of a proper face $C^{\prime}\leq C$. Denote by $\phi_{1},\dots,\phi_{m^{\prime}}$ a basis of covectors orthogonal to $C^{\prime}$. Consider covectors $\eta_{m^{\prime}+1},\dots,\eta_{n}$ vanishing on $Q$ so that $\phi_{1},\dots,\phi_{m^{\prime}},\eta_{m^{\prime}+1},\dots,\eta_{n}$ form a basis in $Q^{\star}$. We can rewrite the system (2) as
    $(\partial_{\phi_{1}}f)^{\xi}=\dots=(\partial_{\phi_{m^{\prime}}}f)^{\xi}=(\partial_{\eta_{m^{\prime}+1}}f)^{\xi}=\dots=(\partial_{\eta_{n}}f)^{\xi}=0$ (4)
    Note that the subsystem of the last $n-m^{\prime}$ equations coincides with the following:
    $(\partial_{\eta_{m^{\prime}+1}}f)|_{C^{\prime}}=\dots=(\partial_{\eta_{n}}f)|_{C^{\prime}}=0$ (5)
    By Proposition 5.19 the latter system (5) has finitely many solutions in the factor-torus $(\mathbb{C}\setminus 0)^{n}/\langle z^{\phi_{1}},\dots,z^{\phi_{m^{\prime}}}\rangle$, where $\langle z^{\phi_{1}},\dots,z^{\phi_{m^{\prime}}}\rangle$ is the subtorus generated by the corresponding one-parameter subgroups (formally speaking, $\phi_{i}$ are affine functionals and here we take only their linear part, but this minor abuse of notations simplifies the formula).
    Consider each solution $x_{\alpha}$ of the system (5) in $(\mathbb{C}\setminus 0)^{n}/\langle z^{\phi_{1}},\dots,z^{\phi_{m^{\prime}}}\rangle$ separately. Let us substitute it in the first $m^{\prime}$ equations of system (4). We obtain a system of $m^{\prime}$ equations in the torus $\langle z^{\phi_{1}},\dots,z^{\phi_{m^{\prime}}}\rangle$ with supports contained in $\pi_{C^{\prime}}(\mathbf{P}\setminus C^{\prime})$, where $\pi_{C^{\prime}}:\mathbb{R}^{n}\to\mathbb{R}^{m^{\prime}}$ is the lattice affine projection whose kernel is the affine span $\text{aff}(C^{\prime})$. The restriction $\xi_{\pi}$ of $\xi$ to $\pi_{C^{\prime}}(\mathbb{R}^{n})$ is well-defined since $\xi$ is constant on $C^{\prime}$. Note that the affine span $\text{aff}((\pi_{C^{\prime}}(\mathbf{P}\setminus C^{\prime}))^{\xi_{\pi}})\subset\mathbb{R}^{m^{\prime}}$ does not contain the origin $\pi_{C^{\prime}}(C^{\prime})$, since $C^{\prime}\cap P$ is the maximal face of $P$ on which the maximum of $\xi$ is attained. Thus the same argument as in item 1. (using Proposition 5.19) implies that the system has no solutions. Considering every $x_{\alpha}$ in the finite set of solutions of the system (5) separately, we obtain that the system (4) has no solution for generic $f$ with support $\mathbf{P}$.
@@
Consider a finite lattice set $\mathbf{P}$, its convex hull $P=\operatorname{Conv}(\mathbf{P})$, and a point $Q\in\mathbb{Q}^{n}$. Suppose that for any face $F\in\mathfrak{A}_{\mathbf{A}}(Q)$, the set $\mathbf{P}$ is smooth at $\mathbf{P}\cap F$. Then we have $\nu_{Q}^{e}(\mathbf{A})=\nu_{{C_{Q}(\mathbf{P})}}(P)$, see Remark 5.21.
@@
5.4 Gelfand-Kapranov-Zelevinsky polytopes via $\mathbf{e}$-Newton number
In this subsection we recall Gelfand-Kapranov-Zelevinsky polytopes (GKZ-polytopes in short) from [GKZ94] and show that the $e$-Newton number computes their support function. This subsection is completely trivial from the point of view of GKZ-polytopes theory. It is not needed for our further results, but it is a good illustration of $e$-Newton numbers.
5.4.1 Preliminaries: Gelfand-Kapranov-Zelevinsky polytopes
Consider a finite set $\mathbf{A}\subset\mathbb{Z}^{n}$. Denote by $\mathbb{C}[\mathbf{A}]$ the set of Laurent polynomials with exponents of their monomials contained in $\mathbf{A}$. Sard’s theorem implies that for generic polynomial $f\in\mathbb{C}[\mathbf{A}]$ with support $\mathbf{A}$ the hypersurface $\{f=0\}$ is smooth in the torus $(\mathbb{C}\setminus 0)^{n}$. Consider all polynomials $f\in\mathbb{C}[\mathbf{A}]$ such that the corresponding hypersurface $\{f=0\}\subset(\mathbb{C}\setminus 0)^{n}$ is not smooth. The closure of this set is denoted by $\nabla_{\mathbf{A}}$.
Definition 5.35.
[GKZ94, Definition 1.2 from Chapter 9] If the set $\mathbf{A}\subset\mathbb{Z}^{n}$ has the property that $\nabla_{\mathbf{A}}\subset\mathbb{C}[\mathbf{A}]$ is a subvariety of codimension 1, then by the $\mathbf{A}$-discriminant we mean an irreducible integral polynomial $\mathbf{D}_{\mathbf{A}}(f)$ in the coefficients $a_{\omega},\omega\in\mathbf{A}$ of $f\in\mathbb{C}[\mathbf{A}]$, which vanishes on $\nabla_{\mathbf{A}}$. Such a polynomial is uniquely determined up to sign. If $\text{codim}\nabla_{\mathbf{A}}>1$, we set $\mathbf{D}_{\mathbf{A}}=1$ and say that the set $\mathbf{A}$ is dual defective.
We recall a classification of dual defective sets in §6.1; for more details, see [CD22].
The Newton polytope $N_{\mathbf{A}}$ of $\mathbf{D}_{\mathbf{A}}$ is computed in [GKZ94], its tropical fan is computed in [DFS07]. A similar problem for singular hyperspaces with codimension two singularities ($A_{2}$ or two $A_{1}$) is solved in [Es18, Section 3.5].
5.4.2 Support function of GKZ-polytope via $\mathbf{e}$-Newton number
Definition 5.36.
Consider a lattice polytope $P\in\mathbb{R}^{n}$ and a covector $\xi\in(\mathbb{Z}^{n})^{*}$. The support function $\xi(P)$ is the maximum of the functional $\xi$ on $P$.
Note that from the support function one can recover the whole polytope. For more details on polyhedral geometry see e.g. [Ew].
Remark 5.37.
Consider polynomial $f\in\mathbb{C}[z_{1},\dots,z_{n}]$ with Newton polytope $P\subset\mathbb{R}^{n}$ and a covector $\xi=(\xi_{1},\dots,\xi_{n})$. Then for generic $(a_{1},\dots,a_{n})$ the number of intersections of $\{f=0\}$ with the curve $(a_{1}z_{1}^{\xi_{1}},\dots,a_{n}z_{n}^{\xi_{n}})$ in the complex torus $(\mathbb{C}\setminus 0)^{n}$ (counted with multiplicities) is equal to $\xi(P)+(-\xi)(P)$.
The latter remark implies the following proposition which allows to recover the Newton polytope of $\mathbf{D}_{\mathbf{A}}$ up to a shift (since $\mathbf{D}_{\mathbf{A}}$ is homogeneous by [GKZ94, Chapter 9, Proposition 1.3.]). For $\mathbf{A}\subset\mathbb{Z}^{n}$ we denote by $\langle\mathbf{A}-\mathbf{A}\rangle$ the subgroup of $\mathbb{Z}^{n}$ generated by the set $\{a_{1}-a_{2}:a_{1},a_{2}\in\mathbf{A}\}$ and say that $\mathbf{A}\subset\mathbb{Z}^{n}$ is spanning, if $\langle\mathbf{A}-\mathbf{A}\rangle=\mathbb{Z}^{n}$. In the following proposition we assume $\mathbf{A}$ to be spanning, but it is easy to extend it to non-spanning sets by multiplying the support function by the index of the corresponding sublattice (see [Es10, Remark on p.18]).
Proposition 5.38.
Consider a finite spanning lattice set $\mathbf{A}\subset\mathbb{Z}^{n}$ and a covector $\xi\in(\mathbb{Z}^{|\mathbf{A}|})^{*}$ with non-negative coordinates. Denote by $\mathbf{A}_{+\xi}$ the set in $\mathbb{Z}^{n}\oplus\mathbb{Z}^{1}_{\geq 0}$ consisting of all the points of the form $(\mathbf{a},0),(\mathbf{a},\xi_{\mathbf{a}}),\mathbf{a}\in\mathbf{A}$. Denote by $O_{\infty}$ the infinite point along the ray $\mathbb{Z}^{1}_{\geq 0}$. Then we have:
$\xi(N_{\mathbf{A}})+(-\xi)(N_{\mathbf{A}})=\nu^{e}_{O_{\infty}}(\mathbf{A}_{+\xi})$
Note that the difference of the lattice volume $\operatorname{Vol}(\operatorname{Conv}(\mathbf{A}_{+\xi}))$ and the $e$-Newton number $\nu^{e}_{O_{\infty}}(\mathbf{A}_{+\xi})$ is consistent with the GKZ-decomposition formula [GKZ94, Chapter 10 Theorem 1.2] (in our notations, see [Es10, Proposition 2.10]).
Remark 5.39.
The Newton number $\nu_{O_{\infty}}^{e}(\cdot)$ from Proposition 5.38 also computes the number of complex critical points of the Bloch variety [FS24, §2]. This computation refines [FS24, Theorems A and 3.8].